OpenStudy (anonymous):

(x^2-5)/(3x^3-7x) Is this symmetric about the origin?

ganeshie8 (ganeshie8):

\[\large y = \dfrac{x^2-5}{3x^3-7x}\] like this ?

OpenStudy (anonymous):

Yes

ganeshie8 (ganeshie8):

one way to check is by replacing x and y with -x and -y and see if you get the same equation

ganeshie8 (ganeshie8):

** it is symmetric about the origin if both the points (x, y) and (-x, -y) exist on the graph of the function

OpenStudy (anonymous):

In different words y=f(x) is symmetric about the origin if f(-x)=-f(x)

OpenStudy (anonymous):

Thank you for help